OpenStudy (anonymous):

1 7(3/14) 13 - + ----- = - 2 10 20 I suck at fractions DX

OpenStudy (anonymous):

what am i solving for? there are no variables

OpenStudy (anonymous):

the last problem you answered for me i have to do a check and idk how to fix this part

OpenStudy (anonymous):

wait does that make sense?

OpenStudy (anonymous):

well 7(3/4) = 21/14 = 3/2 so (3/2) / 10 = 3/20 now 1/2 = 10/20 so 10/20 +3/20 = 13/20